OpenStudy (anonymous):

I am having a hard time doing the lewis structure for ClF3? If we do the calculation there should only be 2 bonds for Cl but that clearly is impossible as we have 3 fluorines?

OpenStudy (anonymous):

To find the number of bonds it's number of electrons needed to complete the valence shell - number of valence electrons. I got (5)(8) - 32 = 4 so 4/2=2 bonds but for some reason my prof changed the 2 to a 3?

OpenStudy (anonymous):

\[ClF _{3}\] has \[3(7)+1(7) = 28\] electrons to place around the atoms. We know we need to place 6 electrons around each fluorine to get them 8 electrons and there are at least single bonds left over. \[3(6) + 3(2) = 24\] This leaves us with 4 electrons or 2 electron pairs. This is a trigonal bipyramidal for the electron shape with a t-shaped molecular shape. So your drawing should have a line to each fluorine with 6 dots around each and 2 extra dots on the chlorine. \[3(6) + 3(2) + 2(2) = 28\]